OpenStudy (ray1998):

Write an equation of a parabola with the given vertex and focus. Vertex (4,-1); Focus (4,-2) What is the equation for the parabola? y= Thank you in advance!

OpenStudy (nubeer):

this would really help you out.. if u still dont get it u can ask but its quite explained here.. https://www.algebra.com/algebra/homework/Quadratic-relations-and-conic-sections/Quadratic-relations-and-conic-sections.faq.question.48599.html

OpenStudy (xariazs):

You know (h,k) and you know the focus from which you can find the p. h=4, k=-1 (x-h)^2=4p(y-k) (x-4)^2=4(-1)(y-(-1)) Written more nicely you get (x-4)^2=-4(y+1)

OpenStudy (ray1998):

Thank you guys!